OpenStudy (anonymous):

Simplify the Sum. 5√8+2√72

OpenStudy (anonymous):

\[5\sqrt{8}+2\sqrt{9*8} = 5\sqrt{8}+6\sqrt{8} = 11\sqrt{}\]

OpenStudy (anonymous):

\[11\sqrt{8} ****\]